OpenStudy (anonymous):

what do you do to check whether a number is rational or irrational. In your explanation use an example of an irrational and rational number.

OpenStudy (er.mohd.amir):

rational number are which can be written in form of p/q where p,q are integer like .333333333333333333333333333.............=1/3

OpenStudy (er.mohd.amir):

and irrational can not writeen in p/q form like sqrt2=1.414.............. can not written p/q or pi=3.1415.......... or e=2.71...........

OpenStudy (anonymous):

Than you @Er.Mohd.AMIR

OpenStudy (anonymous):

*thank